find the accumulated value of an investment of 25,000 for 4 years at an interest rate of 5%, if the money is a compounded semiannually; b. coumpounded quaterly; c. compounded monthly d. coumpounded continuously.
Answer #1
I'll try the first one. The remaining cases are similar and serve as exercises for you (as they should).
Principal, P = 25000
Period, t = 6 months.
Time, T = 4 years.
Number of periods, n = 4/0.5=8
Annual interest rate, = 5%
Interest rate per period, i = 5%*0.5=2.5%

Future value (compound interest)
= P(1+i)^n
= 25000*(1+2.5%)^8
= 25000*1.2184
= $30,460.07
Answer #2
1
answered by: Isuckatmath